# Improving Multimodal Learning with Multi-Loss Gradient Modulation

### v1

@@ 1 Introduction
Combining data from several modalities such as vision, text, audio, and time series has significantly improved performance across many tasks and has proven particularly advantageous in cases of noisy or unreliable sources [25, 15, kontras2024core]. However, studies show that the inclusion of a new modality doesn’t always benefit, and can even impair, model performance [24]. As explained by Huang et al. [12], different modalities compete with each other, resulting in underperforming modalities. These modalities exhibit inferior performance in their multimodal-trained encoders compared to their unimodal counterparts, assessing each modality encoder independently post-training, suggesting potential imbalances or inefficiencies in the integrated training process. This finding contradicts the assumption that more information necessarily improves task understanding.
@@ 2.2 Multi-task Learning
The methods that belong in this category target the sub-optimal unimodal encoders by employing dedicated losses for each modality. It has been observed that incorporating unimodal classifiers enhances multimodal learning, as demonstrated in [23, kontras2024core]. Building on this foundation, Wang et al. [24] suggest to dynamically adjust the weights of the unimodal losses based on an overfitting-to-generalization ratio, however accessing that information requires a separate validation set. Du et al. [3] introduce a teacher-student schema with distillation losses for direct guidance of unimodal encoders. The Prototypical Modal Rebalance (PMR) method [4] circumvents the use of additional parameters for classifiers by leveraging prototype-based classifiers and distance-based losses. This approach maintains a similar logic, wherein imbalanced performance is compensated for by the appropriate unimodal loss. However, balancing the unimodal losses alone disregards the effects caused by the multimodal loss.
@@ 2.3 Other Approaches
Additional studies explore further approaches to multimodal learning, not fitting neatly into previously discussed categories. MMCosine [28] preconditions late fusion by standardizing both the feature vectors and the weights dedicated to each modality, equalizing their contribution to the final prediction. Gat et al. [6] explore a regularization technique to enhance each modality’s contribution by maximizing functional entropy, estimated through prediction differences after input perturbations. This method, however, increases the model’s sensitivity to those perturbations. Adaptive Gradient Modulation (AGM) [14] takes the contribution of each modality by utilizing zero-masking Shapley [shapley1953value, 16] values. This allows for balancing on models of any structure, however at the cost of increased computational demands due to multiple forward passes.

### v2

@@ 1 Introduction
Combining data from several modalities such as vision, text, audio, and time series has significantly improved performance across many tasks and has proven particularly advantageous in cases of noisy or unreliable sources [28, 16, 14, 21, 22]. However, studies show that the inclusion of a new modality doesn’t always benefit, and can even impair, model performance [27]. As explained by Huang et al. [12], different modalities compete with each other, resulting in underperforming modalities. These modalities exhibit inferior performance in their multimodal-trained encoders compared to their unimodal counterparts, assessing each modality encoder independently post-training, suggesting potential imbalances or inefficiencies in the integrated training process. This finding contradicts the assumption that more information necessarily improves task understanding.
@@ 2.2 Multi-task Learning
The methods that belong in this category target the sub-optimal unimodal encoders by employing dedicated losses for each modality. It has been observed that incorporating unimodal classifiers enhances multimodal learning, as demonstrated in [26, 14]. Building on this foundation, Wang et al. [27] suggest to dynamically adjust the weights of the unimodal losses based on an overfitting-to-generalization ratio, however accessing that information requires a separate validation set. Du et al. [3] introduce a teacher-student schema with distillation losses for direct guidance of unimodal encoders. The Prototypical Modal Rebalance (PMR) method [4] circumvents the use of additional parameters for classifiers by leveraging prototype-based classifiers and distance-based losses. This approach maintains a similar logic, wherein imbalanced performance is compensated for by the appropriate unimodal loss. However, balancing the unimodal losses alone disregards the effects caused by the multimodal loss.
@@ 2.3 Other Approaches
Additional studies explore further approaches to multimodal learning, not fitting neatly into previously discussed categories. MMCosine [31] preconditions late fusion by standardizing both the feature vectors and the weights dedicated to each modality, equalizing their contribution to the final prediction. Gat et al. [6] explore a regularization technique to enhance each modality’s contribution by maximizing functional entropy, estimated through prediction differences after input perturbations. This method, however, increases the model’s sensitivity to those perturbations. Adaptive Gradient Modulation (AGM) [15] takes the contribution of each modality by utilizing zero-masking Shapley [shapley1953value, 17] values. This allows for balancing on models of any structure, however at the cost of increased computational demands due to multiple forward passes.